A wagon with a mass of 50t moves at a speed =36km/h. By how much does the internal energy of the brakes of the wagon increase at its complete braking, if 60% of the kinetic energy is converted into internal.


1
Expert's answer
2021-03-02T18:04:23-0500

The initial amount of kinetic energy was:


"K_i = \dfrac{mv^2}{2}"

where "m = 50t = 5\times10^4kg" is the mass, and "v = 36km\/h = 10m\/s" is the initial speed. The final kinetic energy is 0 (since the wagon stops completely). If 60% of the kinetic energy is converted into internal, then the result internal energy increase is:


"Q = 0.6K_i\\
Q = 0.6\cdot \dfrac{5\times 10^4\cdot 10^2}{2} = 1.5\times 10^6J = 1.5 MJ"

Answer. 1.5 MJ.